The partners Carl, Sean, and Frank decide to liquidate on August 1, 2018. The partners share profit/losses using 25%, 45%, and 30% respectively. The B/S of the partners is as follows:

Cash $60,000 A/P $130,000
Loan Receivable, Carl $40,000 Loan Payable, Sean $60,000
Other Assets $500,000 Carl, Capital $140,000
Total Assets $600,000 Sean, Capital $160,000
Frank, Capital $110,000
Total Liab & Capital $600,000

The disposal of Other Assets with a carrying amount of $300,000 realized $250,000. Liquidation expenses are expected to total $40,000. Any partner with a negative capital balance is insolvent. Prepare a schedule of the safe cash payments that can be made to each partner.
Answer #1

Assets Sold for a Loss:

The disposal of Other assets are mentioned in the question as $300,000 but they are sold for $250,000.

This is a $50,000 loss to the partnets and will result in reducing both their assets and capital accounts.

Date Description Dr ($) Cr ($)
2018, Aug 1 Cash 250,000
Loss or Gain on Realization 50,000
Other Assest 300,000
(To record the sale of Other Assets)
2018, Aug 1 Carl, Capital 12,500
Sean, Capital 22,500

Frank, Capital

15,000
Loss or Gain on Realization 50,000
(To record the closing of the Loss to Partners' Capital Accounts)
2018, Aug 1 Accounts Payable 130,000
Cash 130,0000
(To Record the payment of the Creditors)

Working Notes : 25%, 45% & 30% for Carl, Sean and Frank respectively sharing loss.

50,000 * 25% = 12,500

50,000*45% = 22,500

50,000*30% = 15,000

Date Description Dr ($) Cr ($)
2018, Aug 1 Carl, Capital 10,000
Sean, Capital 18,000
Frank, Capital 12,000
To Liquidation Expenses 40,000
(To record liquidation Expenses)

Working Notes:

Liquidation Exp 40,000 between 25,45 & 30 % respectively.

40,000*25% = 10,000

40,000*45% = 18,000

40,000*30% = 12,000

Carls, Capital

12,500 140,000
10,000 117,500

Sean Capital

22,500 160,000
60,000 77,500
18,000 59,500

Frank Capital,

15,000 110,000
12,000 83,000

Other Assets

300,000 300,000

Accounts Payable

130,000 130,000

Loss or Gain Realiation

50,000 50,000

Cash

60,000 130,000
250,000
180,000

There is $180,000 left in the Cash Account, distributing the remaining Safe cash payments according to the equities of the partners.

Date Description Dr ($) Cr ($)
2018, Aug 1 Carl, Capital 45,000
Sean, Capital 81,000
Frank, Capital 54,000
Cash 180,000
To record the closing of the Partnership Books

Working Notes:

180,000* 25 % , 45% & 30%

After this entry is posted all the accounts have a zero balance and partnership is terminated.